# Combinatorial Client-Master Multiagent Deep Reinforcement Learning for Task Offloading in Mobile Edge Computing

### v1

@@ Abstract.
Recently, there has been an explosion of mobile applications that perform computationally intensive tasks such as video streaming, data mining, virtual reality, augmented reality, image processing, video processing, face recognition, and online gaming. However, user devices (UDs), such as tablets and smartphones, have a limited ability to perform the computation needs of the tasks. Mobile edge computing (MEC) has emerged as a promising technology to meet the increasing computing demands of UDs. Task offloading in MEC is a strategy that meets the demands of UDs by distributing tasks between UDs and MEC servers. Deep reinforcement learning (DRL) is gaining attention in task-offloading problems because it can adapt to dynamic changes and minimize online computational complexity. However, the various types of continuous and discrete resource constraints on UDs and MEC servers pose challenges to the design of an efficient DRL-based task-offloading strategy. Existing DRL-based task-offloading algorithms focus on the constraints of the UDs, assuming the availability of enough storage resources on the server. Moreover, existing multiagent DRL (MADRL)–based task-offloading algorithms are homogeneous agents and consider homogeneous constraints as a penalty in their reward function. We proposed a novel combinatorial client-master MADRL (CCM_MADRL) algorithm for task offloading in MEC (CCM_MADRL_MEC) that enables UDs to decide their resource requirements and the server to make a combinatorial decision based on the requirements of the UDs. CCM_MADRL_MEC is the first MADRL in task offloading to consider server storage capacity in addition to the constraints in the UDs. By taking advantage of the combinatorial action selection, CCM_MADRL_MEC has shown superior convergence over existing MADDPG and heuristic algorithms.
@@ 1. Introduction
Recently, there has been an explosion of mobile applications that perform computation-intensive tasks, such as video streaming, virtual reality, augmented reality, image processing, video processing, face recognition, and online gaming Zhang et al. (2013); Chen et al. (2023); Birhanu Engidayehu et al. (2022); Kan et al. (2018). However, UDs, such as tablets and smartphones, have limited ability to perform the computation tasks of these applications. Mobile Cloud Computing (MCC) has been considered the key technology to meet the computing needs of UDs by offloading their computation tasks to the cloud Mahenge et al. (2022). One of the challenges of MCC is latency caused by the distance of the MCC server from the UDs Sajnani et al. (2018). MEC has emerged as a promising technology to address the challenges of MCC and the increasing computing demands of UDs by providing the MCC service on the edge of the network.
Task offloading in MEC has become an attractive solution to meet the diverse computing needs of UDs Islam et al. (2021) by distributing computational tasks between UDs and MEC servers. Many existing task-offloading algorithms use traditional convex optimization methods for single-agent offloading scenarios Sadatdiynov et al. (2023). DRL has recently gained attention in task offloading due to its advantage in reducing online computational complexity Nie et al. (2023) and adapting to dynamic changes Liu et al. (2022a). However, the existence of various types of resource constraints on UDs and MEC servers and the combination of discrete, continuous, and combinatorial action spaces pose challenges to the design of an efficient DRL-based task-offloading strategy. UDs have limitations such as finite battery life and limited computational capabilities Kan et al. (2018); Yan et al. (2018), as well as quality of service (QoS) requirements such as latency. Similarly, MEC servers come with storage constraints. DRL techniques, such as deep Q network (DQN), have yielded encouraging results by modeling the task-offloading problem as MDP with deep neural network (DNN) for function approximation Liu et al. (2022b). However, due to the curse of dimensionality, DQN is insufficient for learning with large discrete action spaces Dulac-Arnold et al. (2015) and a combination of continuous and discrete action spaces Zhang et al. (2020). Although MADDPG-based task-offloading algorithms can handle continuous action spaces, the representation of discrete and continuous action spaces still poses a challenge Zhang et al. (2020); Jiang et al. (2023). Despite the advances of MADRL in task offloading, such as cooperative offloading decisions Nguyen et al. (2023) and mixed continuous and discrete action spaces Zhang et al. (2020); Jiang et al. (2023), most existing MADRL-based algorithms still formulate the constraints as a penalty in their reward function.
@@ 4.5. Algorithms
33:     $Append\left(Q^{N}_{i},Q\left(S_{i},A^{new}_{i},S_{i,n},A^{new}_{i,n},\phi\right)\right)$, $\quad\forall n\in N$ with $A_{i,n}^{mas}=1$, and for $i=1$ to $M$
@@
40:     Compute the gradient for the client: $\nabla_{\phi_{n}}J(\phi_{n})\leftarrow-\frac{1}{Length(tarQ)}\mathop{\raisebox{-2.15277pt}{\Huge$\displaystyle\Sigma$}}_{j=1}^{Length\left(tarQ\right)}\nabla_{\phi_{n}}tarQ_{i}$
@@
We used DDQN van2016deep and prioritized experience replay Schaul et al. (2016) in the CCM_MADRL algorithms for better efficiency in the training.
@@ 5.2. Experimental Settings
The experimental setting is provided in Table 2. Note that all UDs have the same minimum battery, power, and resource allocation threshold, but their maximum budget is generated from a uniform distribution. As described in Section 2, the experimental setting is customized from the settings in Nguyen et al. (2023) and in Zhang et al. (2020). The typical storage capacity of modern servers is GBs and TBs. However, because we chose a small experimental setting due to computational resources, we considered a storage constraint of 400 MB so that the task offloading problem is combinatorial to the server. Evaluation episodes are seeded with their index. A discount factor of 0.99 is used in this experiment. We used a seed of 37 for the reproducibility of the simulation environment. The source code is publicly available at https://github.com/TesfayZ/CCM_MADRL_MEC.
@@
$|N|$
50
$P^{max}$
24 dBm
$K$
10
$P^{min}$
1 dBm
$\tau_{n}$
[0.1-0.9] s
$p_{n}^{max}$
[$P^{min}$ $P^{max}$] dBm
$W$
40 MHz
$p_{n}^{min}$
$P^{min}$ dBm
$f_{max}$
1.5 GHz
$f_{min}$
0.4 GHz
$f_{n}^{max}$
[$f^{min}$- $f^{max}$] GHz
$b^{max}$
3.2 MJ
$f_{n}^{min}$
$f^{min}$ GHz
$b^{min}$
0.5 MJ
$g_{n}$
[5-14] dB
$f_{e}$
4 GHz
$z_{n}$
[1-50] MB
$z_{e}$
400 MB
$c_{n}$
[300 - 737.5] cycles
$U_{e}$
8
$\kappa$
$5\times 10^{-27}$
$\lambda_{1},\lambda_{2}$
0.5, 0.5
$e_{n}$
0.001 J
$b_{n}^{max}$
[$b^{min}$- $b^{max}$] MJ
$b_{n}^{min}$
$b^{min}$ MJ
@@ 5.4. Results
Next, to see the impact on the battery level, we repeat the above experiment by changing the number of steps per episode to 100 and the $b_{max}$ to $b_{min}+1J$. Figure 4 (C) shows that CCM_MADRL_MEC has more UDs running below the minimum battery threshold than the benchmark and heuristic algorithms. However, as explained above for Figure 3, the algorithm is trained by computing a scalar reward, which is a sum of energy and time consumption, as seen in Equation (13) but the percentage of UDs that exceed their battery threshold is from the ration of UDs that exceed the battery threshold to the total number of UDs. Therefore, it is affected by the scales given to the deadline penalty and the energy penalty. We used $lambda_{1}$ = $lambda_{2}$ = 0.5 as weight coefficients for processing time and energy consumption. We repeated the algorithm for $lambda_{1}$ = 1 and $lambda_{2}$ = 5 as seen in Figure 5.
Figure 4. Comparison of the CCM_MADRL with the heuristic and MADDPG algorithms on the evaluation environment with learning rates of 0.0001 and 0.001 and $b_{max}$ = $b_{min}+1J$ for the clients and master respectively
We further run the experiment with $lambda_{1}$ = 1 and $lambda_{2}$ = 5 because Figure 4 (B) looks like it is the exact inverse of Figure 4 (B) means $lambda_{1}$ = $lambda_{2}$ = 0.5 has not given enough balance for the energy consumption, which is a very low number, and the time consumption, which is relatively larger. The result is plotted in Figure 5 with a 95% confidence interval of 40 runs, unlike the above two experiments which are plotted with a 95% confidence interval of 10 runs. However, the plots do not show a significant difference with 10 runs and 40 runs. Subplot (B) and subplot (A) are now varied, but subplot (C) shows that CCM_MADRL_MEC has still more UDs running below the minimum battery threshold than the benchmark and heuristic algorithms. Nonetheless, CCM_MADRL_MEC has performed better than the other algorithms in the combined performance and shows more advantage in the number of tasks whose deadlines exceeded. Using scalar reward for multi-objective functions does not represent the underlying problem Vamplew et al. (2022), but it is not within the scope of this work. The work by Hayes et al. (2022) has studied the essence and techniques of multi-objective reinforcement learning.
Figure 5. Comparison of the CCM_MADRL with the heuristic and MADDPG algorithms on the evaluation environment with learning rates of 0.0001 and 0.001 and $b_{max}$ = $b_{min}+1J$ for the clients and master respectively

### v2

@@ Abstract.
Recently, there has been an explosion of mobile applications that perform computationally intensive tasks such as video streaming, data mining, virtual reality, augmented reality, image processing, video processing, face recognition, and online gaming. However, user devices (UDs), such as tablets and smartphones, have a limited ability to perform the computation needs of the tasks. Mobile edge computing (MEC) has emerged as a promising technology to meet the increasing computing demands of UDs. Task offloading in MEC is a strategy that meets the demands of UDs by distributing tasks between UDs and MEC servers. Deep reinforcement learning (DRL) is gaining attention in task-offloading problems because it can adapt to dynamic changes and minimize online computational complexity. However, the various types of continuous and discrete resource constraints on UDs and MEC servers pose challenges to the design of an efficient DRL-based task-offloading strategy. Existing DRL-based task-offloading algorithms focus on the constraints of the UDs, assuming the availability of enough storage resources on the server. Moreover, existing multiagent DRL (MADRL)–based task-offloading algorithms are homogeneous agents and consider homogeneous constraints as a penalty in their reward function. We proposed a novel combinatorial client-master MADRL (CCM_MADRL) algorithm for task offloading in MEC (CCM_MADRL_MEC) that enables UDs to decide their resource requirements and the server to make a combinatorial decision based on the requirements of the UDs. CCM_MADRL_MEC is the first MADRL in task offloading to consider server storage capacity in addition to the constraints in the UDs. By taking advantage of combinatorial action selection, CCM_MADRL_MEC has shown superior convergence over existing benchmark and heuristic algorithms.
@@ 1. Introduction
Recently, there has been an explosion of mobile applications that perform computation-intensive tasks, such as video streaming, virtual reality, augmented reality, image processing, video processing, face recognition, and online gaming Zhang et al. (2013); Chen et al. (2023); Birhanu Engidayehu et al. (2022); Kan et al. (2018). However, UDs, such as tablets and smartphones, have limited ability to perform the computation tasks of these applications. Mobile Cloud Computing (MCC) has been considered the key technology to meet the computing needs of UDs by offloading their computation tasks to the cloud Mahenge et al. (2022). One of the challenges of MCC is latency caused by the distance of the MCC server from the UDs Sajnani et al. (2018). MEC has emerged as a promising technology for addressing the challenges of MCC and the increasing computing demands of UDs by providing MCC services on the edge of the network.
Task offloading in MEC has become an attractive solution to meet the diverse computing needs of UDs Islam et al. (2021) by distributing computational tasks between UDs and MEC servers. Many existing task-offloading algorithms use traditional convex optimization methods for single-agent offloading scenarios Sadatdiynov et al. (2023). DRL is a common solution for task offloading problems due to its advantage in reducing online computational complexity Nie et al. (2023) and adapting to dynamic changes Liu et al. (2022a). However, the existence of various types of resource constraints on UDs and MEC servers and the combination of discrete, continuous, and combinatorial action spaces pose challenges to the design of an efficient DRL-based task-offloading strategy. UDs have limitations such as finite battery life and limited computational capabilities Kan et al. (2018); Yan et al. (2018), as well as quality of service (QoS) requirements such as latency. Similarly, MEC servers come with storage constraints. DRL techniques, such as deep Q network (DQN), have yielded encouraging results by modeling the task-offloading problem as MDP with deep neural network (DNN) for function approximation Liu et al. (2022b). However, due to the curse of dimensionality, DQN is insufficient for learning with large discrete action spaces Dulac-Arnold et al. (2015) and a combination of continuous and discrete action spaces Zhang et al. (2020). Although multiagent deep deterministic policy gradient (MADDPG) algorithms can handle continuous action spaces, the representation of discrete and continuous action spaces still poses a challenge Zhang et al. (2020); Jiang et al. (2023). Despite the advances of MADRL in task offloading, such as cooperative offloading decisions Nguyen et al. (2023) and mixed continuous and discrete action spaces Zhang et al. (2020); Jiang et al. (2023), most existing MADRL-based algorithms still formulate the constraints as a penalty in their reward function.
@@ 4.5. Algorithms
33:     $Append\left(Q^{N}_{i},Q\left(S_{i},A^{new}_{i},S_{i,n},A^{new}_{i,n},\phi\right)\right)$, $\quad\forall n\in N$ with $A^{new}_{i,n}\geq 0$, and for $i=1$ to $M$
@@
40:     Compute the gradient for the client: $\nabla_{\phi_{n}}J(\phi_{n})\leftarrow-\frac{1}{Length(tarQ)}\mathop{\raisebox{-2.15277pt}{\Huge$\displaystyle\Sigma$}}_{j=1}^{Length\left(tarQ\right)}\nabla_{\phi_{n}}tarQ_{j}$
@@
We used DDQN Van Hasselt et al. (2016) and prioritized experience replay Schaul et al. (2016) in the CCM_MADRL algorithms for better efficiency in the training.
@@ 5.2. Experimental Settings
The experimental setting is provided in Table 2. Note that all UDs have the same minimum battery, power, and resource allocation threshold, but their maximum budget is generated from a uniform distribution. As described in Section 2, the experimental setting is customized from the settings in Nguyen et al. (2023) and Zhang et al. (2020). The typical storage capacity of modern servers is GB and TB. However, because we chose a small experimental setting due to computational resources, we considered a storage constraint of 400 MB so that the task offloading problem is combinatorial to the server. Evaluation episodes are seeded with their index. We used a seed of 37 for the reproducibility of the simulation environment. The source code is publicly available at https://github.com/TesfayZ/CCM_MADRL_MEC.
The hyperparameters for the CMMADRL and benchmark algorithms are as follows: a discount factor of 0.99, a minibatch size of 64, a replay memory size of 10,000, and learning rates of 0.0001 for the clients and actors and 0.001 for the master and critic networks. The output activation function is Tanh for the clients and actors, while ReLU is used for the master and critic networks.
The neural network architecture for the clients and actors consists of an input layer with 7 neurons, two hidden layers with 64 and 32 neurons, and an output layer of size 3. The critic network in the benchmark algorithms has an input layer of 500 neurons, representing the combined state and action size for the 50 actors, followed by two hidden layers with 512 and 128 neurons, and a single output neuron. The master agent has similar parameters to the critic, but its input layer has 510 neurons to account for computing the Q-value for each client using a per-action DQN He et al. (2015). All input and hidden layers use linear activation functions.
@@
$|N|$ 50 $P^{max}$ 24 dBm
$K$ 10 $P^{min}$ 1 dBm
$\tau_{n}$ [0.1-0.9] s $p_{n}^{max}$ [$P^{min}$ $P^{max}$] dBm
$W$ 40 MHz $p_{n}^{min}$ $P^{min}$ dBm
$f_{max}$ 1.5 GHz $f_{min}$ 0.4 GHz
$f_{n}^{max}$ [$f^{min}$- $f^{max}$] GHz $b^{max}$ 3.2 MJ
$f_{n}^{min}$ $f^{min}$ GHz $b^{min}$ 0.5 MJ
$g_{n}$ [5-14] dB $f_{e}$ 4 GHz
$z_{n}$ [1-50] MB $z_{e}$ 400 MB
$c_{n}$ [300 - 737.5] cycles $U_{e}$ 8
$\kappa$ $5\times 10^{-27}$ $\lambda_{1},\lambda_{2}$ 0.5, 0.5
$e_{n}$ 0.001 J $b_{n}^{max}$ [$b^{min}$- $b^{max}$] MJ
$b_{n}^{min}$ $b^{min}$ MJ
@@ 5.4. Results
Next, to see the impact on the battery level, we repeat the above experiment by changing the number of steps per episode to 100 and the $b_{max}$ to $b_{min}$+$1J$. Figure 4 (C) shows that CCM_MADRL_MEC has more UDs running below the minimum battery threshold than the benchmark and heuristic algorithms. However, as explained above for Figure 3, the algorithm is trained by computing a scalar reward, which is a sum of energy and time consumption, as seen in Equation (13) but the percentage of UDs that exceed their battery threshold is from the ration of UDs that exceed the battery threshold to the total number of UDs. Therefore, it is affected by the scales given to the deadline penalty and the energy penalty. We used $\lambda_{1}$ = $\lambda_{2}$ = 0.5 as weight coefficients for processing time and energy consumption. We repeated the algorithm for $\lambda_{1}$ = 1 and $\lambda_{2}$ = 5 as seen in Figure 5.
Figure 4. Comparison of the CCM_MADRL with the heuristic and MADDPG algorithms on the evaluation environment with learning rates of 0.0001 and 0.001 for the clients and master respectively, and $b_{max}$ = $b_{min}$+$1J$
We further run the experiment with $\lambda_{1}$ = 1 and $\lambda_{2}$ = 5 because Figure 4 (B) looks like it is the exact inverse of Figure 4 (B) means $\lambda_{1}$ = $\lambda_{2}$ = 0.5 has not given enough balance for the energy consumption, which is a very low number, and the time consumption, which is relatively larger. The result is plotted in Figure 5 with a 95% confidence interval of 40 runs, unlike the above two experiments which are plotted with a 95% confidence interval of 10 runs. However, the plots do not show a significant difference with 10 runs and 40 runs. Subplot (B) and subplot (A) are now varied, but subplot (C) shows that CCM_MADRL_MEC has still more UDs running below the minimum battery threshold than the benchmark and heuristic algorithms. Nonetheless, CCM_MADRL_MEC has performed better than the other algorithms in the combined performance and shows more advantage in the number of tasks whose deadlines exceeded. Using scalar reward for multi-objective functions does not represent the underlying problem Vamplew et al. (2022), but it is not within the scope of this work. The work by Hayes et al. (2022) has studied the essence and techniques of multi-objective reinforcement learning.
Figure 5. Comparison of the CCM_MADRL with the heuristic and MADDPG algorithms on the evaluation environment with learning rates of 0.0001 and 0.001 for the clients and master respectively, $b_{max}$ = $b_{min}$+$1J$, and $\lambda_{2}$ = 5